OpenStudy (anonymous):

Reduce the following rational expression to its lowest terms: 10v^4 + 25 v^4 w^5 OVER 15vy^6

OpenStudy (anonymous):

\[10v^4+25v^4w^5/15vy^6\]\[=v^3(5w^5+2)/3y^6\]\[=5v^3w^5+2v^3/3y^6\]\[=v^3[(5^w5/3y^6)+(2/3y^6)]\]\[=(5v^3w^5/3y^6)+(2v^3/3y^6)\]

OpenStudy (anonymous):

OMG thanxz 4 da steps this is going to help me with the rest of the hw !!!!!!!!!!!!!!!!!!!!!!! ;-)